# RatBodyFormer: Rodent Body Surface from Keypoints

### v1

@@ C.1 Individual-dependent Parameters
Figure F: L2 error histograms of D2-a and D2-b. Each vertical bar shows the mean L2 error of the histogram in the same color. The blue and orange histograms show the errors with and without optimizing the individual-dependent parameters, respectively. Training with individual-dependent parameter optimization consistently improves the estimation accuracy.
@@
Figure G: L2 error histograms of 5- and 14-week-old rats. Each vertical bar indicates the mean L2 error of the histogram in the same color. The blue and orange histograms show the errors with and without optimizing the individual-dependent parameters, respectively. Inference-time optimization of the individual-dependent parameters consistently improves the estimation accuracy.
C.2 Data Normalization
Fig. H shows the effect of our data normalization applied to the input of RatBodyFormer described in L283 of the main text using D1 splits. These results clearly demonstrates that the normalization significantly improves the accuracy.
@@ C.2 Data Normalization
Figure H: L2 error histograms of D1-a, D1-b, D1-c, and D1-d. Each vertical bar shows the mean L2 error of the histogram in the same color. Our data normalization improves the average L2 error by about 3.8 mm to 9.2 mm.

### v3

@@ E.1 Position Consistency Loss
Figure H: L2 error histograms of D2-a and D2-b. Each vertical bar indicates the mean L2 error of the histogram in the same color. The blue and orange histograms show the errors with and without a position consistency loss, respectively. Even with a position consistency loss, the estimation error of own body surface points does not significantly degrade.
E.2 Individual-dependent Parameters
This section evaluates the effect of the individual-dependent parameters, i.e., point-wise scaling parameters $\mathbf{C}$ and translation parameters $\mathbf{T}$. Fig. I shows the effect of optimizing the individual-dependent parameters at inference time using the D3 split in Table 1 of the main text. In this split, RatBodyFormer is trained using 7-, 9-, and 11-week-old rats with manually-annotated and semi-automatically annotated data, and evaluated with 5- and 14-week-old rats using 3D points obtained by LiDAR and shape-from-silhouette as described in the main text. For the results of inference without optimizing individual-dependent parameters, we set $\mathbf{C}=1$ and $\mathbf{T}=0$. We can observe that the optimization improves the estimation accuracy.
@@ E.2 Individual-dependent Parameters
Figure I: L2 error histograms of 5- and 14-week-old rats. Each vertical bar indicates the mean L2 error of the histogram in the same color. The blue and orange histograms show the errors with and without optimizing the individual-dependent parameters, respectively. Optimizing the individual-dependent parameters consistently improves the estimation accuracy.
E.3 Data Normalization
Fig. J shows the effect of our data normalization applied to the input of RatBodyFormer using D1 splits. These results clearly demonstrate that the normalization significantly improves the accuracy.
@@ E.3 Data Normalization
Figure J: L2 error histograms of D1-a, D1-b, D1-c, and D1-d. Each vertical bar shows the mean L2 error of the histogram in the same color. Our data normalization improves the average L2 error by about 5.5 mm to 8.0 mm.
Appendix F Error Propagation from Keypoints
In this section, we demonstrate the sensitivity of RatBodyFormer to the accuracy of keypoint estimation. Fig. K shows L2 error histograms when adding 3 mm and 8 mm noise to input 10 keypoint 3D coordinates at inference time of D1-a in Tab. A. Even when 8 mm noise is added to the input keypoints, the average body surface estimation error increases by no more than 3 mm compared to the case without noise. This shows that the error propagation from keypoints is not significant.
Figure K: L2 error histograms of D1-a. Each vertical bar shows the mean L2 error of the histogram in the same color. The blue, orange and green histograms show the errors without noise, with 3 mm noise, and with 8 mm noise, respectively. Adding 8 mm noise to the input keypoints does not cause the average estimation error to increase by more than 3 mm from the case without noise. Keypoint estimation inaccuracy does not significantly affect the estimation accuracy of the body surface.
Appendix G Initialization in RatBodyFormer
In this section, we demonstrate the impact of initial estimation accuracy on body surface estimation accuracy. Fig. L shows the relationship between the initial body surface estimation error by ARAP deformation and the final estimation error by RatBodyFormer. Even when the initial estimation error by ARAP deformation exceeds 20 mm, the final estimation error by RatBodyFormer is below 10 mm in most cases, indicating low sensitivity to initial estimation errors.
Figure L: The relationship between the initial estimation error by ARAP deformation and the final estimation error by RatBodyFormer. RatBodyFormer’s final estimation error is below 10 mm in most cases even with initial estimation errors by ARAP deformation exceeding 20 mm, which indicates its low sensitivity to initial estimation errors.